OpenStudy (kkutie7):

A commuter train travels between two stations by accelerating at the rate a for a time t1 and then slowing down with a negative acceleration of -5a for a time t2. The two stations are a distance d apart and the train never reaches its maximum speed. The train stops at each station. Find times t1 and t2.

OpenStudy (kkutie7):

I'm also not allowed to use any other equation except these: \[v^{\rightarrow}=\frac{dr^{\rightarrow}}{dt}\] \[a^{\rightarrow}=\frac{dv^{\rightarrow}}{dt}\] \[v^{\rightarrow}=v _{0}^{\rightarrow}+at^{\rightarrow}\] \[r^{\rightarrow}=r _{0}^{\rightarrow}+v _{0}^{\rightarrow}t+\frac{1} {2}a^{\rightarrow}t^{2}\]

OpenStudy (anonymous):

during time t1: r - ro = vo*t1 + (1/2)*a*t1^2 but vo = 0 d1 = (1/2)*a*t1^2 and when t = t1: v = a*t1 during time t2: vo = v, from the previous step r - ro = vo*t2 +(1/2)*(-5a)*t2^2 d2 = a*t1*t2 - (5/2)*a*t2^2 and when t = t2: 0 = a*t1 -5a*t2 t2 = (1/5)*t1 d1+d2 = d d = (1/2)*a*t1^2 + a*t1*t2 - (5/2)*a*t2^2 plug in the relation between t2 and t1: d = (1/2)*a*t1^2 + a*t1*t2 - (5/2)*a*t2^2 d = (1/2)*a*t1^2 + a*(1/5)*t1^2 - (5/2)*a*((1/5)*t1)^2 d = (1/2)*a*t1^2 + a*(1/5)*t1^2 - (1/10)*a*t1^2 (1/2 + 1/5 - 1/10)*a*t1^2 = d (3/5)*a*t1^2 = d t1^2 = (5/3)*d/a t1 = sqrt((5/3)*d/a) Now find t2: t2 = (1/5)*t1 t2 = (1/5)*sqrt((5/3)*d/a)